# Higher-order homogenised riblet boundary conditions

### v1

@@ 5 Homogenisation of the 2D Navier–Stokes equations
Flow past a riblet surface can be considered as having three characteristic lengths: the riblet period $s^{*}$, the distance to the opposing wall $L^{*}$ and a dynamic length which will be taken to be the viscous length $\ell^{*}=\nu^{*}/u^{*}_{\tau}$. For definiteness here $u^{*}_{\tau}$ will be the characteristic perturbation velocity that in a turbulent flow is usually defined from the mean wall-shear stress, but for the purpose of what follows any other reference velocity $u^{*}_{R}$ would do as well, as long as it correctly represents the order of magnitude of velocities, the essential underlying assumption being that all three components of velocity scale proportionally to each other (as is ordinarily the case in a near-wall turbulent flow but not, for instance, in a laminar boundary layer). The ratio $L^{*}/\ell^{*}=L^{*}u^{*}_{\tau}/\nu^{*}$ is one of the possible definitions of Reynolds number, and decides the type of flow. When $L^{*}\ll\ell^{*}$ the Reynolds number is low, $\ell^{*}$ becomes irrelevant, and we are in the Stokes regime dealt with in the previous sections. When $\ell^{*}\ll L^{*}$, the Reynolds number is high and we are in the Navier-Stokes, possibly turbulent, regime. Near the wall $L^{*}$ now becomes irrelevant, and the asymptotic-expansion parameter $\varepsilon$ must be chosen to be the ratio $\varepsilon=s^{*}/\ell^{*}$, the same parameter that is frequently denoted as $s^{+}$ (or riblet period in wall units).
@@
In a nondimensionalization with reference length $\ell^{*}$, reference velocity $u^{*}_{\tau}$, reference time $\ell^{*}/u^{*}_{\tau}$, and reference pressure $\mu^{*}u_{\tau}^{*}/\ell^{*}=\rho^{*}u_{\tau}^{*2}$, all constant dimensional coefficients in these equations disappear and they become
@@
The highest derivatives in this equation are the same as in the transverse Stokes equation (4.2), and the required boundary conditions are also the same. In particular, 2 boundary conditions are required at each of the two walls of a strip-like domain, one of which is textured, and a solid wall is characterized by $\psi=\psi_{\text{w}}={}$constant (where the constant may, on one of the two walls, be assumed to be zero) and $\partial\psi/\partial n=0$.
@@ 5.1 Outer expansion
where the small parameter is, as already discussed, $\varepsilon=s^{*}/\ell^{*}=s^{+}$. Upon inserting (5.4) into (5.3) and collecting equal powers of $\varepsilon$, $\psi_{0}$ still obeys the same equation (5.3), but with no-slip and impermeability conditions on the virtual flat wall, $\psi_{0}(y,0,t)=\psi_{0,z}(y,0,t)=0$. From the (numerical, say) solution of the $\psi_{0}$-equation we can obtain, among other things, the values of $\psi_{0,zz}(y,0,t)$ and $\psi_{0,zzz}(y,0,t)$. With this information available, we can formally reconstruct the solution by Cauchy series. The beginning of this series will be identical to (4.10) up to third order, and only start to differ from the fourth-order term which is derived from the governing differential equation:
@@ 5.2 Inner expansion
(5.7) may then look like $\varPsi_{0}$ shall obey a nonlinear equation. However, it must be remembered that in §4.2 $\varPsi_{0}$ and $\varPsi_{1}$ turned out to be identically zero, and the expansion effectively started with $\varPsi_{2}$. There is a simple qualitative explanation for this behaviour: owing to the boundary conditions $\psi=\psi_{z}=0$, the streamfunction $\psi$ is O$(z^{2})$ near the wall $z=0$, and thus proportional to $\varepsilon^{2}Z^{2}$ in inner coordinates, and of the second order in $\varepsilon$. This estimate is unchanged in the Navier-Stokes problem. With the expansion (5.6) starting to be non-zero from $i=2$, (5.7) produces the following hierarchy:
@@ 5.3 Orders 0, 1, 2
where numerical constants $a_{1}$, $a_{2}$, $b_{2}$ and $c_{2}=-b_{2}$ are defined in (4.17), (4.25), (4.19) respectively, and $a_{2}$ is expected to be zero for left-right-symmetric riblets as will be shown later to be the case.
@@
This is not the most useful form, however. Whereas in (5.9) up to the third $z$-derivative of $\psi$ is allowed to appear, in primitive variables $v_{zz}$ and $w_{z}$ can be substituted from the momentum and continuity equations (5.2), and actually must be substituted in order to maintain compatibility with the Cauchy expansion in the developments to be presented later in the 3D extensions of this argument. (When the continuity equation, for instance, becomes 3-dimensional $u_{x}+v_{y}+w_{z}=0$, $v_{y}$ and $-w_{z}$ become no longer equivalent; $v_{y}$, and not $w_{z}$, is the quantity that appears on par with $u_{x}$ in the Cauchy expansion.)
@@
The first-order condition contains, as already stressed, the classical protrusion height for the wall-parallel velocity and zero for the wall-normal velocity. The wall-parallel second-order condition contains the wall-parallel pressure gradient in a similar manner to the longitudinal condition (3.11); however, just as (3.11) contains an additional second-order term proportional to $u_{yz}$, both of (5.10f) contain an additional term proportional to $v_{yz}$. One of the coefficients of these terms, $a_{2}$, vanishes for left-right-symmetric riblets (and in reality for any kind of riblets, as will be discovered later), but not $-c_{2}$ which always equals $b_{2}$ (see appendix A).
@@ 5.4 Third order
Since the l.h.s. of this equation is still linear, its solution can be obtained by simply adding to the expression (4.30) of $\varPsi_{4}$ the two additional contributions $\overline{\varPsi_{43}}(Y,Z)\,\psi_{0,zzt}(y_{0},0,t)+\overline{\varPsi_{44}}(Y,Z)\,\psi^{2}_{0,zz}(y_{0},0,t)$, where $\overline{\varPsi_{43}}$ obeys the equation
@@
Note: $e_{3}^{(nl)}$, together with $a_{2}$, $b_{3}$ and $c_{3}$, is one of those coefficients which are forbidden by symmetry, and thus zero, for left-right-symmetric riblets, because when $\overline{\varPsi_{21}}$ is even, the r.h.s. of (5.13) is odd-symmetric with respect to $Y$ and so is $\overline{\varPsi_{44}}$. It follows that for left-right-symmetric riblets nonlinearity only intervenes at an even higher than third order in $\varepsilon$.
@@
Let us now expand each of $u,v,w,p$ in a series of powers of $\varepsilon$, using capital letters for the terms of the inner expansion as in $u=\sum\varepsilon^{i}U_{i}$. As suggested by the previous sections, we may expect the inner expansions of $u,v,w$ (first derivatives of $\psi$) to start at order 1 and the inner expansion of pressure (akin to a second derivative of $\psi$) to start at order 0. We thus obtain the following hierarchy.
@@
$\boldsymbol{\cdots\cdots\cdots}$
One can observe as much: i) each order is governed by the same, linear, system of l.h.s. equations, driven in various ways by the solutions of previous orders; ii) this system of equations neatly decouples into a 2D Poisson problem for $U_{n}$ and a steady 2D Stokes problem for $V_{n},W_{n},P_{n-1}$, exactly those problems that were studied in the previous sections of this paper; iii) with respect to the previous sections there are some new forcing terms, most notably in the continuity equation; iv) since the hierarchical problem is linear we can exploit superposition, taking the already acquired results for granted and focusing on each new additional term in turn.
@@ 6.1 First order
Nothing new to say. The first-order equations are homogeneous just as they were in the previous sections, and driven by the matching conditions only. They lead to the longitudinal and transverse protrusion heights of Luchini et al. (1991) according to (2.13) and (4.17).
@@ 6.2 Second order
($p_{0,x}$ being the same as the constant pressure gradient formerly considered in §3), where $\overline{\varPhi_{21}}$ was defined in (2.17), $\overline{U_{22}}$ was defined in (3.9), and $\overline{U_{23}}$ obeys
@@
As to the r.h.s. of (6.4d), it makes a streamfunction impossible to use to derive the velocity field, and the simplest choice is to solve these equations numerically in primitive form. Nonetheless, their solution can be written as the linear superposition of a solution with non-zero forcing from $v_{0,z}(x,y_{0},0,t)$ and zero r.h.s., which is exactly the one already dealt with in §4.2.3, and one with $v_{0,z}(x,y_{0},0,t)=0$ and non-zero r.h.s. of (6.4d), which can be represented through a normalized solution $\overline{V_{23}},\overline{W_{23}},\overline{P_{13}}$ that solves
@@ 6.3 Third order
The procedure should by now be self-evident. Each stage of the $\varepsilon$-expansion entails a growing number of contributions, which superpose linearly and can be handled independently of each other. All of the terms already included in the 2D analysis are present. As compared to the 2D Navier-Stokes equations discussed in §5.4, the third order (6.5) of the 3D Navier-Stokes problem produces the following additions.
@@ 6.3.1 Longitudinal strain
Indeed, the analogy is so precise that in the transverse equivalent boundary condition (5.16) it suffices to replace $v_{0,zt}$ by $v_{0,zt}-v_{0,xxz}$ (and $w_{0,zt}$ by $w_{0,zt}-w_{0,xxz}$ if they were present; but, by continuity and the no-slip conditions on $u_{0}$ and $v_{0}$, $w_{0,zt}=w_{0,xxz}=0$). In the longitudinal momentum equation, where the time derivative was not included in the analysis of §3, this contribution to $U_{3}$ can be written as $\overline{U_{33}}(Y,Z)[u_{0,zt}(x,y_{0},0,t)-u_{0,xxz}(x,y_{0},0,t)]$, where $\overline{U_{33}}$ is the solution of
@@ 6.3.2 Nonlinear terms
can be subdivided into a transverse part, which is identical to the one already studied in §5.4, and a longitudinal part which represents transport of longitudinal momentum caused by the crossflow velocities. The r.h.s. for the longitudinal part can be written as $[\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}_{,Z}]v_{0,z}u_{0,z}$. Therefore the corresponding contribution to $U_{3}$ is $\overline{U_{34}}(Y,Z)v_{0,z}u_{0,z}$, where
@@ 6.3.3 Cross-coupling
To convert (6.18) into (6.19) it mostly suffices to apply Cauchy-Riemann relationships analogous to (6.6) term by term. This is true, in particular, for $\overline{P_{01}}$, already available from (6.6) itself, and for $\overline{P_{12}}$, obtained from a similar set of equations applied to $\overline{\varPsi_{32}}$, whereas $\overline{P_{13}}$ directly comes from (6.11). $\overline{P_{11}}$ requires special care, owing to the presence of the term proportional to $Y$ in the square brackets.
@@
$\overline{P_{11}}$ must be obtained from (6.20), with a boundary condition $\overline{P_{11}}(Y,\infty)\sim-Z$.
The contribution to $U_{3}$ resulting from all the terms of (6.19) together, to be added to (2.28), will be
@@
where $\overline{U_{22}}$ and $\overline{U_{23}}$ were already defined in (3.9) and (6.8) respectively, and $\overline{U_{35}}$, $\overline{U_{36}}$, $\overline{U_{37}}$ obey
@@
Each term in turn must be inserted into four problems similar to (6.11), one with r.h.s. $-\overline{\varPhi_{11}}(Y,Z)$ yielding the same $\overline{V_{23}},\overline{W_{23}},\overline{P_{13}}$ as (6.11), one with r.h.s. $-\overline{\varPhi_{21}}(Y,Z)-Y\overline{\varPhi_{11}}(Y,Z)$ yielding $\overline{V_{35}},\overline{W_{35}},\overline{P_{25}}$, one with r.h.s. $-\overline{U_{22}}(Y,Z)$ yielding $\overline{V_{36}},\overline{W_{36}},\overline{P_{26}}$ and one with r.h.s. $-\overline{U_{23}}(Y,Z)$ yielding $\overline{V_{37}},\overline{W_{37}},\overline{P_{27}}$.
Just as with (6.20), the term $-\overline{\varPhi_{21}}(Y,Z)-Y\overline{\varPhi_{11}}(Y,Z)$ requires special care for the presence of a part proportional to $Y$. Inserting this binomial into a Stokes problem like (6.11) gives
@@
$\left\{\begin{array}[]{l}\varepsilon^{3}f_{3}^{(u_{xyz})}u_{xyz}+\varepsilon^{3}f_{3}^{(p_{xx})}p_{xx}+\varepsilon^{3}f_{3}^{(v_{xxz})}v_{xxz}\\ \varepsilon^{3}g_{3}^{(u_{xyz})}u_{xyz}+\varepsilon^{3}g_{3}^{(p_{xx})}p_{xx}+\varepsilon^{3}g_{3}^{(v_{xxz})}v_{xxz}\\ \end{array}\right.$
@@
$f_{3}^{(u_{xyz})}=\lim_{Z\rightarrow\infty}[\overline{V_{35}}(Y,Z)+Z^{3}/6],\quad f_{3}^{(p_{xx})}=\lim_{Z\rightarrow\infty}\overline{V_{36}}(Y,Z),\quad f_{3}^{(v_{xxz})}=\lim_{Z\rightarrow\infty}\overline{V_{37}}(Y,Z)$ (6.24)
@@
$\begin{split}&g_{3}^{(u_{xyz})}=\lim_{Z\rightarrow\infty}[\overline{W_{35}}(Y,Z)+(h_{2}+f_{2})Z],\quad g_{3}^{(p_{xx})}=\lim_{Z\rightarrow\infty}[\overline{W_{36}}(Y,Z)+Z^{3}/6+h_{2}^{(p_{x})}Z],\\ &g_{3}^{(v_{xxz})}=\lim_{Z\rightarrow\infty}[\overline{W_{37}}(Y,Z)+h_{2}^{(v_{xz})}Z].\end{split}$ (6.25)

### v2

@@ 5 Homogenisation of the 2D Navier–Stokes equations
Flow past a riblet surface can be considered as having three characteristic lengths: the riblet period $s^{*}$, the distance to the opposing wall $L^{*}$ and a dynamic length which will be taken to be the viscous length $\ell^{*}=\nu^{*}/u^{*}_{\tau}$. For definiteness here $u^{*}_{\tau}$ will be the characteristic perturbation velocity that in a turbulent flow is usually defined from the mean wall-shear stress, but for the purpose of what follows any other reference velocity $u^{*}_{R}$ would do as well, as long as it correctly represents the order of magnitude of near-wall velocities. Another defining feature of the distinguished limit that we adopt here is that all three components of velocity scale proportionally to each other, i.e. share the same reference velocity $u^{*}_{R}=u^{*}_{\tau}$. This is ordinarily the case in a near-wall turbulent flow but not, for instance, in a laminar boundary layer. The ratio $L^{*}/\ell^{*}=L^{*}u^{*}_{\tau}/\nu^{*}$ is one of the possible definitions of Reynolds number, and decides the type of flow. When $L^{*}\ll\ell^{*}$ the Reynolds number is low, $\ell^{*}$ becomes irrelevant, and we are in the Stokes regime dealt with in the previous sections. When $\ell^{*}\ll L^{*}$, the Reynolds number is high and we are in the Navier-Stokes, possibly turbulent, regime. Near the wall $L^{*}$ now becomes irrelevant, and the asymptotic-expansion parameter $\varepsilon$ must be chosen to be the ratio $\varepsilon=s^{*}/\ell^{*}$, the same parameter that is frequently denoted as $s^{+}$ (or riblet period in wall units). This discussion only shows, one could argue, that $\varepsilon$ must be proportional to $s^{+}$, but in the argument of a power expansion a fixed multiplying coefficient (say, $\varepsilon=s^{*}/\ell^{*}$ versus $\varepsilon=s^{*}/(20\ell^{*})$) makes no difference.
@@
In a nondimensionalization with reference length $\ell^{*}$, reference velocity $u^{*}_{\tau}$, reference time $\ell^{*}/u^{*}_{\tau}$, and reference pressure $\mu^{*}u_{\tau}^{*}/\ell^{*}=\rho^{*}u_{\tau}^{*2}$, all constant dimensional coefficients in (5.1) disappear and these equations become
@@
The highest derivatives in this equation are the same as in the transverse Stokes equation (4.2), and the required boundary conditions are also the same. In particular, two boundary conditions are required at each of the two walls of a strip-like domain, one of which is textured, and a solid wall is characterized by $\psi=\psi_{\text{w}}={}$constant (where the constant may, on one of the two walls, be assumed to be zero) and $\partial\psi/\partial n=0$.
@@ 5.1 Outer expansion
where the small parameter is, as already discussed, $\varepsilon=s^{*}/\ell^{*}=s^{+}$. Upon inserting (5.4) into (5.3) and collecting equal powers of $\varepsilon$, $\psi_{0}$ still obeys the same equation (5.3), but with no-slip and impermeability conditions on the virtual flat wall, $\psi_{0}(y,0,t)=\psi_{0,z}(y,0,t)=0$ (as explained near (4.2)). From the (numerical, say) solution of the $\psi_{0}$-equation we can obtain, among other things, the values of $\psi_{0,zz}(y,0,t)$ and $\psi_{0,zzz}(y,0,t)$. With this information available, we can formally reconstruct the solution by Cauchy series. The beginning of this series will be identical to (4.11) up to third order, and only start to differ in the fourth-order term which is derived from the governing differential equation:
@@ 5.2 Inner expansion
[…]
@@ 5.3 Orders 0, 1, 2
where numerical constants $a_{1}$, $a_{2}$, $b_{2}$ and $c_{2}=-b_{2}$ are defined in (4.18), (4.26), (4.20) respectively, and $a_{2}$ is expected to be zero for left-right-symmetric riblets (and in fact for any riblets, as will later be shown to be the case).
@@
This is not the most useful form, however. Whereas in (5.9) up to the third $z$-derivative of $\psi$ is allowed to appear, in primitive variables $v_{zz}$ and $w_{z}$ can be substituted from the momentum and continuity equations (5.2), and actually must be substituted in order to maintain compatibility with the Cauchy expansion in the developments to be presented later in the 3D extensions of this argument. (When the continuity equation, for instance, becomes 3-dimensional as $u_{x}+v_{y}+w_{z}=0$, $v_{y}$ and $w_{z}$ become no longer interchangeable; $u_{x},v_{y}$, without $w_{z}$, is the pair that appears in the Cauchy expansion and formulas derived from it.)
@@
The first-order condition contains, as already stressed, the classical transverse protrusion height $a_{1}$ for the wall-parallel velocity and zero for the wall-normal velocity. The wall-parallel second-order condition contains the wall-parallel pressure gradient in a similar manner to the longitudinal condition (3.11); however, just as (3.11) contains an additional second-order term proportional to $u_{yz}$, both of (5.10f) contain an additional term proportional to $v_{yz}$. One of the coefficients of these terms, $a_{2}$, vanishes for left-right-symmetric riblets (and in reality for any kind of riblets, as will be discovered later), but not $c_{2}$ which always equals $b_{2}$ (see appendix A).
@@ 5.4 Third order
Since the l.h.s. of this equation is still linear, its solution can be obtained by simply adding to the expression (4.31) of $\varPsi_{4}$ the two new contributions $\overline{\varPsi_{43}}(Y,Z)\,\psi_{0,zzt}(y_{0},0,t)+\overline{\varPsi_{44}}(Y,Z)\,\psi^{2}_{0,zz}(y_{0},0,t)$ (with an incremental second subscript according to the convention set forth before), where $\overline{\varPsi_{43}}$ obeys the equation
@@
Note: $e_{3}^{(nl)}$, together with $a_{2}$, $b_{3}$ and $c_{3}$, is one of those coefficients which are forbidden by symmetry, and thus zero, for left-right-symmetric riblets, because when $\overline{\varPsi_{21}}$ is even, the r.h.s. of (5.13) is odd-symmetric in $Y$ and so is $\overline{\varPsi_{44}}$. It follows that for left-right-symmetric riblets nonlinearity only intervenes at an even higher than third order in $\varepsilon$.
@@
Let us now expand each of $u,v,w,p$ in a series of powers of $\varepsilon$, using capital letters for the terms of the inner expansion as in $u=\sum\varepsilon^{i}U_{i}$. As suggested by the previous sections, we may expect the inner expansions of $u,v,w$ (first derivatives of $\psi$) to start at order 1 and the inner expansion of pressure (akin to a velocity gradient, or to a second derivative of $\psi$) to start at order 0. We thus obtain the following hierarchy.
@@
[…]
One can observe as much: i) each order is governed by the same, linear, system of l.h.s. equations, driven in various ways by the solutions of previous orders; ii) this system of equations neatly decouples into a 2D Poisson problem for $U_{n}$ and a steady 2D Stokes problem for $V_{n},W_{n},P_{n-1}$, exactly those problems that were solved in the previous sections of this paper; iii) with respect to the previous sections there are some new forcing terms, including those in the continuity equation; iv) just as in §5, nonlinear terms make their first appearance at third order (and not at second order as might have been inferred from their quadratic nature); v) since the hierarchical problem is linear we can exploit superposition, taking the already acquired solutions of the homogeneous problem for granted and adding in the contribution of each new r.h.s. term in turn.
@@ 6.1 First order
Nothing new to add. The first-order equations are homogeneous just as they were in the previous sections, and driven by the matching conditions only. They lead to the longitudinal and transverse protrusion heights of Luchini et al. (1991) according to (2.13) and (4.18).
@@ 6.2 Second order
($p_{0,x}$ being the same as the constant pressure gradient formerly considered in §3), where $\overline{\varPhi_{21}}$ was defined in (2.17), $\overline{U_{22}}$ was defined in (3.9), and the new $\overline{U_{23}}$ obeys
@@
As to the r.h.s. of (6.4d), it makes a streamfunction impossible to use to derive the velocity field, and the simplest choice is to solve (6.4bdd) numerically in primitive form. Nonetheless, their solution can be written as the linear superposition of a solution with non-zero forcing from $v_{0,z}(x,y_{0},0,t)$ and zero r.h.s., which can be derived from a streamfunction and is exactly the one already dealt with in §4.2.3, and a solution with $v_{0,z}(x,y_{0},0,t)=0$ and non-zero r.h.s. of (6.4d), which can be represented through a new normalized triplet $\overline{V_{23}},\overline{W_{23}},\overline{P_{13}}$ that solves
@@ 6.3 Third order
The procedure should by now be self-evident. Each stage of the $\varepsilon$-expansion entails a growing number of contributions, which superpose linearly and can be handled independently of each other. All of the terms already included in the 2D analysis are present. As compared to the 2D Navier-Stokes equations discussed in §5.4, the third order (6.5) of the 3D Navier-Stokes problem produces the following additions, which we shall label according to the convention already adopted in the previous sections, a first digit denoting order and a second sequential digit.
@@ 6.3.1 Longitudinal strain
Indeed, the analogy is so precise that in the transverse equivalent boundary condition (5.15) it suffices to replace $v_{0,zt}$ by $v_{0,zt}-v_{0,xxz}$ (and $w_{0,zt}$ by $w_{0,zt}-w_{0,xxz}$ if they were present; but, by continuity and the no-slip conditions on $u_{0}$ and $v_{0}$, $w_{0,zt}=w_{0,xxz}=0$). In the longitudinal momentum equation, where the time derivative was not included in the analysis of §3, this contribution to $U_{3}$ can be written as $\overline{U_{33}}(Y,Z)[u_{0,zt}(x,y_{0},0,t)-u_{0,xxz}(x,y_{0},0,t)]$, where the new function $\overline{U_{33}}$ is the solution of
@@ 6.3.2 Nonlinear terms
can be subdivided into a transverse part, which is identical to the one already studied in §5.4, and a longitudinal part which represents transport of longitudinal momentum caused by the crossflow velocities. The r.h.s. for the longitudinal part can be written as $\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}_{,Z}]v_{0,z}u_{0,z}$. Therefore the corresponding contribution to $U_{3}$ is a new function $\overline{U_{34}}(Y,Z)v_{0,z}u_{0,z}$, where
@@ 6.3.3 Cross-coupling
To convert (6.18) into (6.19) it mostly suffices to apply Cauchy-Riemann relationships analogous to (6.6) term by term. This is true, in particular, for $\overline{P_{01}}$, already available from (6.6) itself, and for $\overline{P_{12}}$, obtained from a similar set of equations applied to $\overline{\varPsi_{32}}$, whereas $\overline{P_{13}}$ directly comes from (6.11). $\overline{P_{11}}$ requires special care, owing to the presence of the term proportional to $Y$ in the square brackets, as follows.
@@
$\overline{P_{11}}$ must then be obtained from (6.20), with a boundary condition $\overline{P_{11}}(Y,\infty)\sim-Z$.
The contribution to $U_{3}$ resulting from all the terms of (6.19) together, to be added to (2.28), can be written as
@@
where $\overline{U_{22}}$ and $\overline{U_{23}}$ were already defined in (3.9) and (6.8) respectively, and the new functions $\overline{U_{35}}$, $\overline{U_{36}}$, $\overline{U_{37}}$ obey
@@
Each term in turn must be inserted into four problems similar to (6.11), one with r.h.s. $\overline{\varPhi_{11}}(Y,Z)$ yielding the same $\overline{V_{23}},\overline{W_{23}},\overline{P_{13}}$ as (6.11), one with r.h.s. $\overline{\varPhi_{21}}(Y,Z)-Y\overline{\varPhi_{11}}(Y,Z)$ yielding a new triplet $\overline{V_{35}},\overline{W_{35}},\overline{P_{25}}$, one with r.h.s. $\overline{U_{22}}(Y,Z)$ yielding a new triplet $\overline{V_{36}},\overline{W_{36}},\overline{P_{26}}$ and one with r.h.s. $\overline{U_{23}}(Y,Z)$ yielding a new triplet $\overline{V_{37}},\overline{W_{37}},\overline{P_{27}}$.
Just as with (6.20), the term $\overline{\varPhi_{21}}(Y,Z)-Y\overline{\varPhi_{11}}(Y,Z)$ requires special care for the presence of a part proportional to $Y$. Inserting this binomial into a Stokes problem like (6.11) gives
@@
[…]
@@
$\begin{split}&f_{3}^{(u_{xyz})}=\lim_{Z\rightarrow\infty}[\overline{V_{35}}(Y,Z)+Z^{3}/6],\quad f_{3}^{(p_{xx})}=\lim_{Z\rightarrow\infty}\overline{V_{36}}(Y,Z),\\ &f_{3}^{(v_{xxz})}=\lim_{Z\rightarrow\infty}\overline{V_{37}}(Y,Z)\end{split}$ (6.24)
@@
$\begin{split}&g_{3}^{(u_{xyz})}=\lim_{Z\rightarrow\infty}[\overline{W_{35}}(Y,Z)+(h_{2}+f_{2})Z],\\ &g_{3}^{(p_{xx})}=\lim_{Z\rightarrow\infty}[\overline{W_{36}}(Y,Z)+Z^{3}/6+h_{2}^{(p_{x})}Z],\\ &g_{3}^{(v_{xxz})}=\lim_{Z\rightarrow\infty}[\overline{W_{37}}(Y,Z)+h_{2}^{(v_{xz})}Z].\end{split}$ (6.25)
6.4 Fourth and higher order
Although we shall not in this paper pursue the detailed calculation of fourth- and higher-order contributions, the road should now be clear. The next order in the expansion of (6.2), for instance, is
Fourth order:
$U_{4,YY}+U_{4,ZZ}=P_{2,x}-U_{2,xx}+U_{2,t}+U_{1}U_{1,x}+V_{2}U_{1,Y}+V_{1}U_{2,Y}+W_{2}U_{1,Z}+W_{1}U_{2,Z},$ (6.26a)
$V_{4,YY}+V_{4,ZZ}-P_{3,Y}=-V_{2,xx}+V_{2,t}+U_{1}V_{1,x}+V_{2}V_{1,Y}+V_{1}V_{2,Y}+W_{2}V_{1,Z}+W_{1}V_{2,Z},$ (6.26b)
$W_{4,YY}+W_{4,ZZ}-P_{3,Z}=-W_{2,xx}+W_{2,t}+U_{1}W_{1,x}+V_{2}W_{1,Y}+V_{1}W_{2,Y}+W_{2}W_{1,Z}+W_{1}W_{2,Z},$ (6.26c)
$V_{4,Y}+W_{4,Z}=-U_{3,x}.$ (6.26d)
As can be seen, linear terms follow a similar pattern at every order, whereas nonlinear, advective terms produce a convolution of previous orders stemming from their product structure. When these systems are solved in a cascade they will produce a general $n^{\text{th}}$ order equivalent boundary condition for the velocity vector $(u,v,w)$ at the equivalent wall $z=0$ akin to (4.34)-(4.35). Just as (4.34)-(4.35) contained higher and higher lateral derivatives of $\psi_{zz}$ and $\psi_{zzz}$, this boundary condition will contain partial derivatives of order up to n1n-1 with respect to $x,y$ and $t$ of the triplet $(u_{z},v_{z},p)$. As can be imagined the number of terms goes up rapidly with order, though not as rapidly as it might, because many of them will have zero coefficients. Rather than trying to provide a tensorial notation for the general formula, which would be cumbersome, we prefer to name each relevant term individually in the next section.